OpenStudy (anonymous):

For the functions f(x) = 5x − 3 and g(x) = x + 4, which composition produces the greatest output? A.f(g(x)) produces the greatest output. B.g(f(x)) produces the greatest output. C. Both compositions produce the same output. D.Neither composition produces an output.

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

@sweetburger

OpenStudy (anonymous):

f(g(x)) = 5(x + 4) - 3 = 5x + 17 g(f(x)) = (5x - 3) + 4 = 5x + 1 Which is greater, 5x + 17 or 5x + 1?